OpenStudy (anonymous):

okay question is eliminate the parameter and obtain the standard form of the rectangular equation...x=2t-3 and y=3t+3 i have what i think is the answer is it right???

OpenStudy (anonymous):

solve for t in the first equation and substitute that into the y=3t+3... what did you get?

OpenStudy (anonymous):

i got t=y/3-1

OpenStudy (anonymous):

i mean what's your equation in xy after you eliminated the parameter t...

OpenStudy (anonymous):

im not sure wat ur trying to ask:(

OpenStudy (anonymous):

those are parametric equations of x and y.... the question asks to eliminate the parameter t so that your equation is only in terms of x and y... no "t"...

OpenStudy (anonymous):

o. okay then i guess i am not sure how to carry out this problem then...thats my problem

OpenStudy (campbell_st):

let \[t = \frac{(x +3)}{2} \] substitute it into y \[y = \frac{3(x +3)}{2} + 3\] simplifying \[y = \frac{3x}{2} + \frac{9}{2} + 3\] \[y = \frac{3x}{2} + \frac{15}{2}\]

OpenStudy (anonymous):

The whole deal is linear. Let t=0 to get a point (x_0,y_0). Then let t=1 to get (x_1,y_1). Those two points will allow you to get the equation of a line in the Cartesian plane.

OpenStudy (anonymous):

okay can i ask what EXACTLY does eliminating the parameter mean?

OpenStudy (anonymous):

nice one^^^ ... that's another way....:)

OpenStudy (anonymous):

or imply?

OpenStudy (campbell_st):

it means to get rid of the t in both equations... this can only be done by rewriting one equation so that t is the subject. when this is done, substitute into the other equation to get the equation of your line

OpenStudy (anonymous):

okay wat do i substitute?... sorry :(

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

help?

OpenStudy (campbell_st):

ok... if x = 2t - 3 can you make t the subject of the equation..?

OpenStudy (anonymous):

yeah. i think. u mean like let it be t=? right or am wrong?

OpenStudy (campbell_st):

thats what you want... what will t be equal to..?

OpenStudy (anonymous):

t will be = to y/3-1.....?

OpenStudy (campbell_st):

no lets work on x = 2t - 3

OpenStudy (anonymous):

o. okay sorry it'll be x/2+3/2

OpenStudy (campbell_st):

yeap or without the expansion \[t =\frac{ (x +3)}{2}\] now substitute it into \[y = 3t + 3\] this will give an equation with 2 unknowns... x and y

OpenStudy (anonymous):

okay let me plug it in

OpenStudy (anonymous):

wait so the equation i have t=x/2+3/2 that is eliminating the parameter? or is it wen i get the 2 unknowns?

OpenStudy (campbell_st):

a little care with the algebra expansion its \[y = 3 \times \frac{x +3}{2} + 3\]

OpenStudy (anonymous):

wait where did u plug in t=x/2+3/2 just now?

OpenStudy (campbell_st):

I have write it as \[\frac{x + 3}{2} = \frac{x}{2} + \frac{3}{2}\]

OpenStudy (campbell_st):

and remember its 3*x and 3* 3

OpenStudy (anonymous):

okay but y are they = to each other is wat i dont understand

OpenStudy (anonymous):